Memory Sizing

NOTE:  K (kilo) means 1024, not 1000. A byte (B) is 8 bits. A kilobyte (KB) is therefore 8 x 1024 = 8192 bits.

a)  A 32 KB (kilobytes) memory has a 16 bit wordsize. How many words total can be stored in this memory?

_________words

b) A 256 KB memory has a 32 bit wordsize. How many bits are required to address this memory?

_________ bits

c) A computer memory has a 128 bit wordsize. It is made up of 4096 addressable locations. How large is thememory in bytes?

__________ bytes

d) A computer memory has 10 address bits, and its word size is 16 bits. How large is the memory in bytes?

___________bytes
Answer #1

answer 1.)

since word size is 16 bit

32K memory locations hence we have 32*1024 locations=32768=thirty two thousand seven hundred sixty-eight.

answer 2.)

256 KB memory has 256*1024=262144 Bytes

now we have a word size of 32 bit hence we have 4 Bytes per location

hence we will have 65536 locations.

in order for us to address these locations 2n=65535 (n=16)

hence, we need 16 address lines/ bit

c.)

the word size of 128 bit.

4096 address locations.

hence we have (1024*4)4K*16(128/8) bytes memory

64KB of memory.

d.)

number of address lines =16

total address locations=210=1024=1K

since each location has 16 bit that is 2 bytes per location.

1K*2=2KB of memory